# Amortized Active Generation of Pareto Sets

### v2

@@ C.2 Sequence variational distributions
For details on the decoder-transformer with a causal mask see [30, Algorithm 10 & Algorithm 14] for maximum likelihood training and sampling implementation details respectively.
@@
Here $\mathrm{ETransformer}$ is an encoder-transformer, see [30, Algorithm 9] for details, and $\mathrm{NN}_{\phi_{o}}$ is a NN decoder, with a convolutional residual layer for capturing additional local structure. The token generation model is,
@@
$[\mathbf{p}_{0},\ldots,\mathbf{p}_{M}]$ $=\mathrm{ETransformer}_{\phi_{e}}\!(\mathbf{x}^{\prime}[\mathbf{o}\leftarrow x_{\text{mask}}]).$
The same encoder-transformer is shared between the mask and token heads, and only the masked positions are allowed to sample new tokens. For our experiments we bias the categorical distribution so that the original token is not resampled. The choice of the set of seed sequences, $\{\mathbf{x}^{\prime}_{i}\}^{J}_{j=1}$, can drastically impact performance of A-GPS, VSD and CbAS. A simple heuristic that we find works well in practice is to uniformly sample with replacement from the current active Pareto set, $\mathcal{S}_{\textrm{Pareto}}^{t}$, so $\mathbf{x}^{\prime}\sim\mathcal{U}({\mathcal{S}_{\textrm{Pareto}}^{t}})$. This is similar to the strategy used in LaMBO-2, though since they use EHVI, they can maintain only the top-$B$ sequences per round ranked by EHVI, where $B$ is batch size.
Finally, if we use this model as a prior, we drop the conditional mask model, $q_{\phi_{oe}}({\mathbf{o}}|{\mathbf{x}^{\prime}})$, and draw mask-positions independently with a Bernoulli distribution, $\mathrm{Bern}\!\left({o_{m}}\middle|{p_{\text{mask}}=0.15}\right)$. Then we use [30, Algorithm 12] for training.
@@
As mentioned in the text, for the conditional generative models, $q_{\phi}({\mathbf{x}}|{\mathbf{u}})$, for A-GPS, we use the same architectures already discussed, but also learn a sequence prefix embedding from $\mathbf{u}$, as well as a simple 1-hidden layer MLPs for implementing FilM [29] adaptation of the sequence token embeddings,
$\mathbf{e}_{m}=\mathbf{e}_{m-1}\circ(1+f_{\alpha}(\mathbf{u}))+f_{\beta}(\mathbf{u}),\quad\text{where}\quad\mathbf{e}_{0}=f_{\text{prefix}}(\mathbf{u}).$ (37)
Here $\circ$ indicates element-wise product, and $f_{\alpha}$ and $f_{\beta}$ are the FiLM MLPs, and $f_{\text{prefix}}$ is the prefix embedding MLP (often we find just a linear projection is adequate), and $\mathbf{e}_{m}$ are the transformer embeddings. We initialize the transformer weights in these models from their non-conditional counterparts when they are used as priors.
@@ D.1 Synthetic test functions
GMM ($D=2,L=2$): Here each objective is implemented as a Gaussian mixture model, and so is highly multimodal [9]. This is run without observation noise.
@@
A detailed description of these functions and/or their Pareto-front geometries can be found in [13, 41, 4]. We use BoTorch [3] for the implementations of all the synthetic test functions and the baseline MOBO methods. All experiments were initialized using Latin hyper-cube sampling. The original design space is $\mathcal{X}=[0,1]^{D}$ for all problems, and we optimize directly in this space with all methods. The GP based methods use an optimizer that respects these bounds directly, and we clamp A-GPS’s generative model to these bounds. We found transforming the space (e.g. using logit-sigmoid transforms) generally led to worse performance.
@@
The Ehrlich vs. naturalness score benchmark was implemented using the poli benchmarking library [17], where we implemented our own ProtGPT2-based naturalness black box, and used the inbuilt Ehrlich function [33] (not the holo version). For A-GPS we use the aforementioned generative and discriminative models, otherwise the settings are given in Table 6. We use a modified version of the LaMBO-2 algorithm [18] from poli-baselines [17]. We use the following Ehrlich function configurations:
@@ D.3 Bi-grams
For the bi-grams experiment we implemented our own black-box for the poli library based on the experiment in [34]. All architectural details are presented in Appendix C, with additional experimental settings in Table 6. We also report runtimes in Table 7.
@@
For the stability vs. SASA experiment we used bi-objective black-box from the poli library, and the seed sequences and settings from the experiment in [34]. All architectural details are presented in Appendix C, with additional experimental settings in Table 6. We also report runtimes in Table 7, and show some samples from the A-GPS generative model, as well as the empirical Pareto front in Figure 8.
@@ Appendix E Ablation Studies
We introduce a new gradient estimator in Equation 20 based on off-policy importance weighting approximations to the on-policy gradient estimator used by [35]. To test its efficacy, we re-run the Ehrlich vs. naturalness score experiments with this new estimator and the original on-policy variant. We report performance and runtimes in Table 9 for $M=32$. There does not seem to be a consistent difference between the two gradient estimators in terms of hypervolume performance, but runtime is significantly lower for the off-policy estimator, being almost an order of magnitude less for the off-policy variant.
@@ E.2 Off-policy gradient estimator samples
We test the effect on performance of the number of samples, $S$, used for estimating the gradients of the off-policy estimator, Equation 20, used for A-GPS and VSD in Table 10. We also do the same for the CbAS estimator. We generally find that more samples lead to more performance for all methods, however this effect plateaus for A-GPS and VSD starting at $S=256$, whereas CbAS still sees improvement beyond $S=512$. This is also something noted by the original authors in [5].

### v3

@@ C.2 Sequence variational distributions
For details on the decoder-transformer with a causal mask see [phuong2022formal, Algorithm 10 & Algorithm 14] for maximum likelihood training and sampling implementation details respectively.
@@
Here $\mathrm{ETransformer}$ is an encoder-transformer, see [phuong2022formal, Algorithm 9] for details, and $\mathrm{NN}_{\phi_{o}}$ is a NN decoder, with a convolutional residual layer for capturing additional local structure. The token generation model is,
@@
$[\mathbf{p}_{1},\ldots,\mathbf{p}_{M}]$ $=\mathbf{softmax}\!\left({\mathrm{ETransformer}_{\phi_{e}}\!(\mathbf{x}^{\prime}[\mathbf{o}\leftarrow x_{\text{mask}}])}\right).$
Here $\mathbf{softmax}:\real^{|\mathcal{V}|\times M}\to[0,1]^{|\mathcal{V}|\times M}$. The same encoder-transformer is shared between the mask and token heads, and only the masked positions are allowed to sample new tokens. For our experiments we bias the categorical distribution so that the original token is not resampled. The choice of the set of seed sequences, $\{\mathbf{x}^{\prime}_{i}\}^{J}_{j=1}$, can drastically impact performance of A-GPS, VSD and CbAS. A simple heuristic that we find works well in practice is to uniformly sample with replacement from the current active Pareto set, $\mathcal{S}_{\textrm{Pareto}}^{t}$, so $\mathbf{x}^{\prime}\sim\mathcal{U}({\mathcal{S}_{\textrm{Pareto}}^{t}})$. This is similar to the strategy used in LaMBO-2, though since they use EHVI, they can maintain only the top-$B$ sequences per round ranked by EHVI, where $B$ is batch size.
Finally, if we use this model as a prior, we drop the conditional mask model, $q_{\phi_{oe}}({\mathbf{o}}|{\mathbf{x}^{\prime}})$, and draw mask-positions independently with a Bernoulli distribution, $\mathrm{Bern}\!\left({o_{m}}\middle|{p_{\text{mask}}=0.15}\right)$. Then we use [phuong2022formal, Algorithm 12] for training.
@@
As mentioned in the text, for the conditional generative models, $q_{\phi}({\mathbf{x}}|{\mathbf{u}})$, for A-GPS, we use the same architectures already discussed, but also learn a sequence prefix embedding from $\mathbf{u}$, as well as a simple 1-hidden layer MLPs for implementing FilM [perez2018film] adaptation of the sequence token embeddings,
$\mathbf{e}_{m}=\mathbf{e}^{\prime}_{m}\circ(1+f_{\alpha}(\mathbf{u}))+f_{\beta}(\mathbf{u}),\quad\text{where}\quad\mathbf{e}_{0}=f_{\text{prefix}}(\mathbf{u}).$ (37)
Where $\mathbf{e}^{\prime}_{m}$ are the unmodified token embeddings (i.e. outputs of $\mathrm{DTransformer}$ or $\mathrm{ETransformer}$), and $\circ$ indicates element-wise product. $f_{\alpha}$ and $f_{\beta}$ are the FiLM MLPs, and $f_{\text{prefix}}$ is the prefix embedding MLP input into the transformers (often we find just a linear projection is adequate). We initialize the transformer weights in these models from their non-conditional counterparts when they are used as priors.
@@ D.1 Synthetic test functions
GMM ($D=2,L=2$): Here each objective is implemented as a Gaussian mixture model, and so is highly multimodal [daulton2022robust]. This is run without observation noise.
@@
A detailed description of these functions and/or their Pareto-front geometries can be found in [deb2002scalable, zhang2007moea, belakaria2019max]. We use BoTorch [balandat2020botorch] for the implementations of all the synthetic test functions and the baseline MOBO methods. All experiments were initialized using Latin hyper-cube sampling. The original design space is $\mathcal{X}=[0,1]^{D}$ for all problems, and we optimize directly in this space with all methods. The GP based methods use an optimizer that respects these bounds directly, and we clamp A-GPS’s generative model to these bounds. We found transforming the space (e.g. using logit-sigmoid transforms) generally led to worse performance.
@@
The Ehrlich vs. naturalness score benchmark was implemented using the poli benchmarking library [gonzalez-duque2024poli], where we implemented our own ProtGPT2-based naturalness black box, and used the inbuilt Ehrlich function [stanton2024closed] (not the holo version). For A-GPS we use the aforementioned generative and discriminative models, otherwise the settings are given in Table 6. We use a modified version of the LaMBO-2 algorithm [gruver2023protein] from poli-baselines [gonzalez-duque2024poli]. We use the following Ehrlich function configurations:
@@ D.3 Bi-grams
For the bi-grams experiment we implemented our own black-box for the poli library based on the experiment in [stanton2022accelerating]. All architectural details are presented in Appendix C, with additional experimental settings in Table 6. We also report runtimes in Table 7.
@@
For the stability vs. SASA experiment we used bi-objective black-box from the poli library, and the seed sequences and settings from the experiment in [stanton2022accelerating]. All architectural details are presented in Appendix C, with additional experimental settings in Table 6. We also report runtimes in Table 7, and show some samples from the A-GPS generative model, as well as the empirical Pareto front in Figure 8.
@@ Appendix E Ablation Studies
We introduce a new gradient estimator in Equation 20 based on off-policy importance weighting approximations to the on-policy gradient estimator used by [steinberg2025variational]. To test its efficacy, we re-run the Ehrlich vs. naturalness score experiments with this new estimator and the original on-policy variant. We report performance and runtimes in Table 9 for $M=32$. There does not seem to be a consistent difference between the two gradient estimators in terms of hypervolume performance, but runtime is significantly lower for the off-policy estimator, being almost an order of magnitude less for the off-policy variant.
@@ E.2 Off-policy gradient estimator samples
We test the effect on performance of the number of samples, $S$, used for estimating the gradients of the off-policy estimator, Equation 20, used for A-GPS and VSD in Table 10. We also do the same for the CbAS estimator. We generally find that more samples lead to more performance for all methods, however this effect plateaus for A-GPS and VSD starting at $S=256$, whereas CbAS still sees improvement beyond $S=512$. This is also something noted by the original authors in [brookes2019conditioning].
@@ E.3 Prior regularization
NeurIPS Paper Checklist
  1. 1.
    Claims
  3. Answer: [Yes]
  5. 2.
    Limitations
  7. Answer: [Yes]
  9. 3.
    Theory assumptions and proofs
  11. Answer: [Yes]
  13. 4.
    Experimental result reproducibility
  15. Answer: [Yes]
  17. 5.
    Open access to data and code
  19. Answer: [Yes]
  21. 6.
    Experimental setting/details
  23. Answer: [Yes]
  25. 7.
    Experiment statistical significance
  27. Answer: [Yes]
  29. 8.
    Experiments compute resources
  31. Answer: [Yes]
  33. 9.
    Code of ethics
  35. Answer: [Yes]
  37. 10.
    Broader impacts
  39. Answer: [Yes]
  41. 11.
    Safeguards
  43. Answer: [N/A]
  45. 12.
    Licenses for existing assets
  47. Answer: [Yes]
  49. 13.
    New assets
  51. Answer: [Yes]
  53. 14.
    Crowdsourcing and research with human subjects
  55. Answer: [N/A]
  57. 15.
  59. Answer: [N/A]
  61. 16.
    Declaration of LLM usage
  63. Answer: [Yes]